What should u be in order to use u-substitution her | e2x+3 dx ?
Answer #1

) سال بلد Let I=/e² J e 24+3 dx substitute U= 2x+3) du= a da dx= i I = Jeu due) - ļ Seu du = f(eu) to * I = £ 62x+3)+C, whereso, u should be 2x+3.